The demand and supply equations for a certain item are given by
D = –5p + 40
S = –p2 + 30p – 8
Find the equilibrium price

Answer by stanbon(57400)

The demand and supply equations for a certain item are given by
D = –5p + 40
S = –p2 + 30p – 8
Find the equilibrium price
------
Equilibrium occurs when supply equals demand.
-p^2+30p-8 = -5p+40
-p^2+35p-48=0
p^2-35p+48=0
Use the quadratic formula to get:
p = [35+-sqrt(35^2-4*48]/2
p = [35+-sqrt(1033)]/2
p = [35+-32.14]/2
p = $1.43 or p =$33.57
-------
Cheers,
Stan H.